OpenStudy (anonymous):

Which integral will require Integration by Parts more than once? int_{?}^{?} lnxdx int_{?}^{?} xsinxdx int_{?}^{?} (xe^x)dx int_{?}^{?} x^2 sinxdx

OpenStudy (amistre64):

well if I do it; all of them :)

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

You dont have to do all of them! lol

OpenStudy (anonymous):

I thought there was like a format that lets you know. Lol. But i'l just go by whatt imrah told me (:

OpenStudy (amistre64):

i recall a reiterative process from a textbook, but the details elude me ...

OpenStudy (anonymous):

fourth not second sorry

OpenStudy (anonymous):

I take back my answer, it was the last one

OpenStudy (anonymous):

most of these cyclic problem come in form of \[\int{e^x sin(x)dx}\]

OpenStudy (amistre64):

my table is off :)

OpenStudy (amistre64):

\begin{array}c &&dv\\ ---&---&---\\ +&u& v\\ -&du&\int v\\ +&d^2u&\int^2 v\\ -&d^3u&\int^3 v\\ \end{array} that might be better

OpenStudy (amistre64):

\[\int udv=uv-\int vdu+\int^2v\ d^2u -\int^3v\ d^3u+...\]

OpenStudy (anonymous):

so x^2*sinxdx

OpenStudy (amistre64):

-----sin(x) x^2. -cos(x) -2x .-sin(x) +2. cos(x) -0. sin(x) -x^2 cos(x) + 2x sin(x) +2 cos(x) ... sure

OpenStudy (anonymous):

awesome! thanks

OpenStudy (amistre64):

--- sinx x. -cos(x) -1. -sin(x) 0. cos(x) int(x sin(x)) = -x cos(x) + sin(x) ; which is only a single "by parts"

OpenStudy (anonymous):

Good Rule to remember when doing integration by parts is I L A TE --> Inverse Trig, Logarithmic, Algebraic, Trig, Exponential

OpenStudy (amistre64):

is that for u or v :)

OpenStudy (anonymous):

yeah the order which to assign u and v